OpenStudy (anonymous):

Given: 2x – 3(x – 5 + x) = 2x – 3 Prove: x = 3 What are the missing justification steps, in order? Distributive Property, Associative Property of Addition, Combine Like Terms Subtraction Property of Equality, Associative Property of Addition, Addition Property of Equality Subtraction Property of Equality, Commutative Property of Addition, Addition Property of Equality Distributive Property, Commutative Property of Addition, Combine Like Terms

OpenStudy (anonymous):

Mathematical Statement Justification 2x – 3(x - 5 + x) = 2x - 3 Given 2x – 3(x + x – 5) = 2x – 3 Commutative Property of Addition 2x – 3(2x – 5) = 2x – 3 Combine Like Terms 2x – 6x + 15 = 2x – 3 (2x – 6x) + 15 = 2x – 3 – 4x + 15 = 2x – 3 – 6x + 15 = – 3 Subtraction Property of Equality – 6x = – 18 Subtraction Property of Equality x = 3 Division Property of Equality

jimthompson5910 (jim_thompson5910):

going from 2x – 3(2x – 5) = 2x – 3 to 2x – 6x + 15 = 2x – 3 what property are we using here?

OpenStudy (anonymous):

Distributive ?

jimthompson5910 (jim_thompson5910):

yes you got it

jimthompson5910 (jim_thompson5910):

so the answer is either Distributive Property, Associative Property of Addition, Combine Like Terms OR Distributive Property, Commutative Property of Addition, Combine Like Terms

jimthompson5910 (jim_thompson5910):

since those are the only two that start with "distributive property"

jimthompson5910 (jim_thompson5910):

going from 2x – 6x + 15 = 2x – 3 to (2x – 6x) + 15 = 2x – 3 what is happening (which property are we using now)?

OpenStudy (anonymous):

then associative property and then combine like terms

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome